# Engagement-Driven Content Generation with Large Language Models

### v4

@@ Abstract
Keywords Large Language Models  $\cdot$ Engagement Maximization  $\cdot$ Reinforcement Learning
1 Introduction
Large Language Models (LLMs) have recently gained a great deal of attention, especially for their capabilities in one-to-one interactions. Researchers have deeply investigated their skills and limits, comparing LLMs to humans in one-on-one interactions such as, e.g., their abilities in gaming settings (Jacob et al., 2024; Fontana et al., 2024). A recent research direction empirically investigates the susceptibility (Chen et al., 2024) and persuasive capabilities of LLMs (Breum et al., 2024; Salvi et al., 2024), showing that contextual information in specific domains can enhance LLMs’ persuasiveness (Matz et al., 2024). In particular, LLMs have shown competence in generating persuasive messages in public health (Karinshak et al., 2023) and personalized advertising (Meguellati et al., 2024).
@@ 1 Introduction
Figure 1: A visual representation of the proposed framework: (a) A query $q$ is prompted to the LLM, which generates content $t$. (b) The content $t$ is injected into the social network (i.e., posted by a node that corresponds to the LLM agent) and starts to propagate into the network according to an engagement model. (c) The number of active users at the end of the propagation is returned to the LLM as the observed reward. The process repeats until convergence.
@@
2 Background and Related Work
@@
LLM-Agents in complex environments. The integration of Agent-Based Models (ABMs) and Large Language Models has been an emerging area of research. Notable studies, such as those by Vezhnevets et al. (2023) and Gürcan (2024), focus on utilizing LLMs as social agents within ABMs to simulate micro-level behaviors and observe macro-level patterns. These micro-level behaviors can correspond to opinion dynamics models Chuang et al. (2024), and the resulting macro-level patterns can lead to known universal laws such as scale-free networks De Marzo et al. (2023). Recent studies have started to place LLM agents into more complex environments to assess their ability to mimic human behavior Park et al. (2023), and to understand their capabilities in negotiation Bianchi et al. ([n. d.]), code development Qian et al. (2024), and even in tasks like building houses Chen et al. (2023). A concurrent line of inquiry is whether LLMs can learn to align their behavior with social norms Fränken et al. ([n. d.]).
@@
Social media environments. The adoption of information diffusion models in data mining methodologies has been thoroughly investigated for designing viral marketing strategies Kempe et al. (2003); Leskovec et al. (2007); Chen et al. (2010); Lu et al. (2013); Barbieri et al. (2013); Barbieri and Bonchi (2014); Tu and Neumann (2022). Various optimization problems have been tackled such as selecting influential seeds Kempe et al. (2003); Tu and Neumann (2022), designing product features Barbieri and Bonchi (2014), or other variations Lu et al. (2013); Aslay et al. (2015). Yet, a crucial practical dimension remains largely unexamined: how to compose textual content that effectively propagates through a network. In this work, we bridge this gap by proposing a methodology that enables LLMs to generate content maximizing engagement on social platforms.
Proxy for human feedback. Regardless of the type of objectives that machine learning models have, a correct reward function is necessary. In particular, the challenges associated with human feedback in evaluating LLMs have led to alternative approaches, such as LLM-as-a-Judge Huang et al. (2024); Verga et al. (2024) and human-in-the-loop methods Boubdir et al. (2023); Amirizaniani et al. (2024), which aim to enhance scalability and efficiency. While our work shares the goal of reducing the reliance on human experts for data annotation, it introduces a fundamentally novel approach by simulating context-dependent feedback that incorporates the social network structure. Hence, this innovation also minimizes the need for slow, costly, and potentially unethical online experiments, offering a more efficient and ethically sound alternative. To the best of our knowledge, no existing work leverages simulated social media engagement models as feedback mechanisms to fine-tune LLMs, addressing a critical gap in current research.
3 Engagement Model
As discussed in Section 1, we adopt a Reinforcement Learning from Simulated Feedback approach. This method not only helps us avoiding the complexity of live experiments but also allows us to control contextual settings, such as the distribution of opinions and the LLM’s position within the social network. Given our goal to study the capabilities of LLMs in generating content that resonates across the network, we propose a novel engagement model which borrows from information propagation Kempe et al. (2003) the cascade mechanism, and from opinion dynamics Deffuant et al. (2000), the single engagement mechanism.
@@ 3 Engagement Model
Consistently with the Independent Cascade model Kempe et al. (2003), the propagation of a piece of content $t$ is modeled in discrete timesteps and independently of other content that might be propagating concurrently. More precisely, a new content $t$, having a leaning $s_{t}\in[0,1]$ w.r.t. a given topic, is injected in the network (i.e., posted by a node $u\in\mathcal{V}$) at time step $T_{0}$. When a user posts a content, their followers become aware of the content and might engage with it. If they do, they become “active” on $t$ at the subsequent timestep and their followers, in turn, become aware of $t$. The decision of engaging with the content is modeled according to the Bounded Confidence Model Deffuant et al. (2000), which is one of the most studied models in the opinion dynamics literature: a node $u\in V$ engages with the content $t$, to which it is exposed, iff $|s_{t}-x_{u}|<\epsilon$, where $\epsilon\in[0,1]$ is the given confidence bound. This process repeats until no new nodes become active, with each node activating at most once on the specific content $t$.
@@
  3. 3.
    Two black-box functions, $\mathcal{S}$ and $\mathcal{F}$, which, given a content $t$, provide (i) the leaning $s_{t}$ of $t$ with respect to the given topic;
  4. 4.
    A fluency score $f_{t}$ representing the readability of $t$ in natural language; (ii) an engagement model $\mathcal{M}_{\epsilon}$, which determines content propagation based on an associated content value and a bounded confidence level $\epsilon\in[0,1]$.
@@ 4 Proposed Method
For our purposes, we fine-tune the underlying LLM agent using a Reinforcement Learning (RL) mechanism (widely adopted for optimizing LLMs Ouyang et al. (2024)). Generally, RL is suitable for scenarios where an agent needs to learn to interact within a dynamic environment. The agent’s goal is to learn a policy—a strategy or behavioral pattern—through feedback it receives from the environment, in the form of rewards or punishments.
@@
Researchers have refined the above optimization problem to avoid learning instability. In particular, the Proximal Policy Optimization (PPO) Schulman et al. (2017) strategy, which we adopt in our framework, introduces constraints on policy updates through a clipped objective function. This ensures that updates do not destabilize the learning process by limiting the extent of policy changes, thereby providing stability and efficiency in fine-tuning large language models.
@@ 9       $j\leftarrow j+1$
  2. 2.
    Sentiment inference. We use a pre-trained model $\mathcal{S}$ to compute a sentiment value $s_{t}$ associated with the content $t$, i.e., $\mathcal{S}(t)=s_{t}$ (line 4). This utility model is a plug-and-play component that can be adapted to different tasks. In our experiments, we adopt a DistilBERT model Sanh et al. (2019) pretrained for sentiment classification as our function $\mathcal{S}$.
  3. 3.
    Fluency score. To ensure that $t$ conveys proper semantics, we compute a readability value associated with its content. While perplexity is a common measure for evaluating the generation quality of LLMs Gonen et al. (2023), it is sensitive to vocabulary size and sentence length. Studies Miaschi et al. (2020) highlight its low correlation with readability. Martinc et al. (2021), study the adequacy of several readability scores and assess their performance on several datasets. For our purposes, we adopt as our fluency function $\mathcal{F}$ the Flesch–Kincaid (FK) Kincaid et al. (1975) formula (line 4). This score quantifies the grade level required to understand an English statement: the higher the value, the higher its complexity and syntactic quality. It is a crucial component of the final reward $\mathcal{R}$, as it ensures the LLM generates meaningful and fluent statements.
  4. 4.
@@ 5 Experiments Setup
Synthetic data. To rigorously evaluate our proposed methodology across diverse initial conditions, we devise a synthetic data generator, based on random network model inspired by Cinus et al. (2022), and capable of producing realistic networks while allowing to control the key parameters influencing information diffusion, such as homophily, modularity, and opinion distribution. The generator takes as input five parameters – homophily level $\eta$, network modularity $\mu$, shape parameters $\alpha$ and $\beta$ for opinion distribution, the number of nodes $N$ – and outputs a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, also assigning an opinion value $x_{u}\in[0,1]$ to each node $u\in\mathcal{V}$. The generation process starts by constructing the set of edges $\mathcal{E}$ and assigning nodes to non-overlapping communities $c:\mathcal{V}\rightarrow C$ using the LFR model Lancichinetti et al. (2008), known for producing community sizes that adhere to a realistic power-law distribution. The modularity parameter $\mu$ is an input of the LFR model and controls the community segregation within the network. Each community $c_{i}\in C$ gets assigned an initial opinion drawn from a Beta distribution [^3] Beta distribution has been chosen for its flexibility to accommodate various distribution shapes such as uniform, unimodal, and bimodal. $o_{c_{i}}\sim Beta(\alpha,\beta)$. For each node $u\in\mathcal{V}$, its opinion $x_{u}$ is determined by a Bernoulli trial with probability $\eta$ (i.e., homophily). If successful, the node adopts the opinion of its community $x_{u}=o_{c(u)}$; otherwise, it independently samples an opinion from the same Beta distribution.
@@
Figure 2: Different opinion settings (from left to right): positive/negative/neutral/uniform/uniform. All networks except for the last one exhibit high modularity and high homophily (i.e., neighbors have similar opinions). The last graph is indeed an example of network with high modularity but low homophily (i.e., neighbors’ opinions are heterogeneous). Colors represent node opinions, whose values are denoted by the color bar.
@@
Figure 3: Networks with low modularity (left) and high modularity (right). The former presents high connectivity among the nodes, while the latter exhibits clustered communities.
@@
Figure 4: Visualization of the $\mathbb{X}$ follow networks. Nodes are colored on a scale from blue (0: ”Remain”/”No”) to red (1: ”Leave”/”Yes”), and their size resembles their out-degree.
@@
Real data. We also consider two real-world datasets containing content about a controversial topics propagated over $\mathbb{X}$ social network. The first dataset, Brexit, focuses on the remain-leave discourse before the 2016 UK Referendum on exiting the EU Zhu et al. (2020). We use the version in Minici et al. (2022) consisting of 7,589 users, 532,459 directed follow relationships, and 19,963 tweets, each associated with a binary stance. The dataset is preprocessed according to Minici et al. (2022) to assign each user a scalar value $x_{u}\in[0,1]$, referred to as opinion, representing the average stance of the tweets retweeted by user $u$. The stance of each tweet is either 0 (“Remain”) or 1 (“Leave”). The second dataset, Italian Referendum, was gathered during the Italian constitutional referendum in 2016 Lai et al. (2018) and processed according to Minici et al. (2022). This dataset includes 2,894 users, 161,888 edges, and 41,001 tweets. Each user is associated with a scalar value $x_{u}\in[0,1]$ , corresponding to their stance on the referendum question, where values represent support for “No” or “Yes”. The structure of the networks and users’ opinions are depicted in Figure 4. Both datasets exhibit two separated, homophilic, and polarized communities, and a smaller, more heterophilic mixed-opinions community.
@@
Settings. Although other modeling choices are possible [^4] For instance, the Generation from Scratch task: i.e., “Generate a post about [TOPIC]”., in our experiments we focus on a Query Completion Task, as we found that Query Completion achieves faster convergence. We defer a detailed study on the effects of task specification to future work. Specifically, in the synthetic setting, we use the query ($q=$Cats are the most…”), while in the real experiments we adopt $q=$“[Brexit]/[2016 Italian constitutional referendum] is the most…”. Given the query $q$, we obtain $t$ by concatenating $q$ and the generated text $g=\textrm{LLM}_{\theta}(q)$, i.e., $t=q\mathbin{\|}g$. To compute the opinion value $s_{t}$, we adopt a DistilBERT model Sanh et al. (2019), pretrained for sentiment classification.
@@
Concerning the LLM agent, we adopt the 2B version of Gemma Team and et. al. (2024), a lightweight models family released by Google and based on Gemini [^5] https://deepmind.google/technologies/gemini/ technology. Despite its smaller size, Gemma outperforms other models of similar size in tasks such as understanding, reasoning, and safety. We fine-tune the model exploiting the PPOTrainer [^6] https://huggingface.co/docs/trl/main/en/ppo_trainer\#trl.PPOTrainer class from the trl package, which enables training language models with custom rewards. The maximum number of training steps is set to $80$ for the synthetic data and $500$ for the real-world datasets, while the threshold $\tau$ for monitoring the KL-divergence to $75$. We also conducted experiments with other state-of-the-art lightweight LLMs, such as Mistral-7B Jiang et al. (2023), LLaMA2-7B Touvron et al. (2023), and GPT-2 Radford et al. (2019). However, these experiments demonstrated that Gemma-2B provides the best trade-off between performance and computational efficiency. Finally, for the engagement model, we set $\epsilon=0.2$.
@@
Figure 5: Engagement $|\mathcal{A}|$ at each step of our fine-tuning procedure. Columns refer to different opinion distributions: positive/negative/neutral/uniform (left to right). Each plot depicts the trend varying the network structure in terms of modularity and homophily. Colors indicate different positions of the LLM agent. Dashed lines represent the maximum engagement within that configuration, whereas dotted lines indicate its lower-bound (i.e., $|\mathcal{A}|$ produced by the non-finetuned LLM).
@@
Figure 6: Sentiment content $s_{t}$ produced at each step of our fine-tuning procedure. Each column refers to a different opinion distribution: positive/negative/neutral/uniform (from left to right). Each plot depicts the trend varying the network structure in terms of modularity and homophily. Colors indicate different positions of the LLM agent.
Baselines. We evaluate the effectiveness of our approach to fine-tuning the LLM, comparing it with the following baselines: BERT [^7] https://huggingface.co/google-bert/ Kenton and Toutanova (2019), GPT-2 Radford et al. (2019), LLaMA3.1-70B-Instruct [^8] https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct, and ChatGPT-4o [^9] https://openai.com/index/hello-gpt-4o/, in addition to the vanilla (not fine-tuned) Gemma-2B model Team and et. al. (2024). We invoke each model in inference to complete the same query $q$ used in the synthetic experiments (“Cats are the most”) and in the real-world scenarios (“[Brexit]/[2016 Italian constitutional referendum] is the most”). When using BERT, we start from $q$ and masking the next word up to a final length of $10$. For all the other baselines, we set a maximum number of output tokens equal to $128$ and a temperature of $0.7$. Consistently with the fine-tuning process, we retrieve a number of responses equal to the batch size (8), and we further inject them into the network starting from the same source nodes. We average the results in terms of produced engagement.

### v5

@@ Abstract.
journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3736932isbn: 979-8-4007-1454-2/2025/08ccs: Networks Social media networksccs: Human-centered computing Collaborative and social computing systems and toolsccs: Computing methodologies Neural networksccs: Computing methodologies Modeling and simulationccs: Computing methodologies Natural language generation
1. Introduction
Large Language Models (LLMs) have recently gained a great deal of attention, especially for their capabilities in one-to-one interactions. Researchers have deeply investigated their skills and limits, comparing LLMs to humans in one-on-one interactions such as, e.g., their abilities in gaming settings (Jacob et al., 2024; Fontana et al., 2024). A recent research direction empirically investigates the susceptibility (Chen et al., [n. d.]) and persuasive capabilities of LLMs (Breum et al., 2024; Salvi et al., 2024), showing that contextual information in specific domains can enhance LLMs’ persuasiveness (Matz et al., 2024). In particular, LLMs have shown competence in generating persuasive messages in public health (Karinshak et al., 2023) and personalized advertising (Meguellati et al., 2024).
@@ 1. Introduction
Figure 1. A visual representation of the proposed framework: (a) A query $q$ is prompted to the LLM, which generates content $t$. (b) The content $t$ is injected into the social network (i.e., posted by a node that corresponds to the LLM agent) and starts to propagate into the network according to an engagement model. (c) The number of active users at the end of the propagation is returned to the LLM as the observed reward. The process repeats until convergence.
@@
2. Background and Related Work
@@
LLM-Agents in complex environments. The integration of Agent-Based Models (ABMs) and Large Language Models has been an emerging area of research. Notable studies, such as those by Vezhnevets et al. (2023) and Gürcan (2024), focus on utilizing LLMs as social agents within ABMs to simulate micro-level behaviors and observe macro-level patterns. These micro-level behaviors can correspond to opinion dynamics models (Chuang et al., 2024), and the resulting macro-level patterns can lead to known universal laws such as scale-free networks (De Marzo et al., 2023). Recent studies have started to place LLM agents into more complex environments to assess their ability to mimic human behavior (Park et al., 2023), and to understand their capabilities in negotiation (Bianchi et al., [n. d.]), code development (Qian et al., 2024), and even in tasks like building houses (Chen et al., 2023). A concurrent line of inquiry is whether LLMs can learn to align their behavior with social norms (Fränken et al., [n. d.]).
@@
Social media environments. The adoption of information diffusion models in data mining methodologies has been thoroughly investigated for designing viral marketing strategies (Kempe et al., 2003; Leskovec et al., 2007; Chen et al., 2010; Lu et al., 2013; Barbieri et al., 2013; Barbieri and Bonchi, 2014; Tu and Neumann, 2022). Various optimization problems have been tackled such as selecting influential seeds (Kempe et al., 2003; Tu and Neumann, 2022), designing product features (Barbieri and Bonchi, 2014), or other variations (Lu et al., 2013; Aslay et al., 2015). Yet, a crucial practical dimension remains largely unexamined: how to compose textual content that effectively propagates through a network. In this work, we bridge this gap by proposing a methodology that enables LLMs to generate content maximizing engagement on social platforms.
Proxy for human feedback. Regardless of the type of objectives that machine learning models have, a correct reward function is necessary. In particular, the challenges associated with human feedback in evaluating LLMs have led to alternative approaches, such as LLM-as-a-Judge (Huang et al., 2024; Verga et al., 2024) and human-in-the-loop methods (Boubdir et al., 2023; Amirizaniani et al., 2024), which aim to enhance scalability and efficiency. While our work shares the goal of reducing the reliance on human experts for data annotation, it introduces a fundamentally novel approach by simulating context-dependent feedback that incorporates the social network structure. Hence, this innovation also minimizes the need for slow, costly, and potentially unethical online experiments, offering a more efficient and ethically sound alternative. To the best of our knowledge, no existing work leverages simulated social media engagement models as feedback mechanisms to fine-tune LLMs, addressing a critical gap in current research.
3. Engagement Model
As discussed in Section 1, we adopt a Reinforcement Learning from Simulated Feedback approach. This method not only helps us avoiding the complexity of live experiments but also allows us to control contextual settings, such as the distribution of opinions and the LLM’s position within the social network. Given our goal to study the capabilities of LLMs in generating content that resonates across the network, we propose a novel engagement model which borrows from information propagation (Kempe et al., 2003) the cascade mechanism, and from opinion dynamics (Deffuant et al., 2000), the single engagement mechanism.
@@ 3. Engagement Model
Consistently with the Independent Cascade model (Kempe et al., 2003), the propagation of a piece of content $t$ is modeled in discrete timesteps and independently of other content that might be propagating concurrently. More precisely, a new content $t$, having a leaning $s_{t}\in[0,1]$ w.r.t. a given topic, is injected in the network (i.e., posted by a node $u\in\mathcal{V}$) at time step $T_{0}$. When a user posts a content, their followers become aware of the content and might engage with it. If they do, they become “active” on $t$ at the subsequent timestep and their followers, in turn, become aware of $t$. The decision of engaging with the content is modeled according to the Bounded Confidence Model (Deffuant et al., 2000), which is one of the most studied models in the opinion dynamics literature: a node $u\in V$ engages with the content $t$, to which it is exposed, iff $|s_{t}-x_{u}|<\epsilon$, where $\epsilon\in[0,1]$ is the given confidence bound. This process repeats until no new nodes become active, with each node activating at most once on the specific content $t$.
@@
  3. (3)
    Two black-box functions, $\mathcal{S}$ and $\mathcal{F}$, which, given a content $t$, provide (i) the leaning $s_{t}$ of $t$ with respect to the given topic; (ii) A fluency score $f_{t}$ representing the readability of $t$ in natural language .
  4. (4)
    An engagement model $\mathcal{M}_{\epsilon}$, which determines content propagation based on an associated content value and a bounded confidence level $\epsilon\in[0,1]$.
@@ 4. Proposed Method
For our purposes, we fine-tune the underlying LLM agent using a Reinforcement Learning (RL) mechanism (widely adopted for optimizing LLMs (Ouyang et al., 2024)). Generally, RL is suitable for scenarios where an agent needs to learn to interact within a dynamic environment. The agent’s goal is to learn a policy—a strategy or behavioral pattern—through feedback it receives from the environment, in the form of rewards or punishments.
@@
Researchers have refined the above optimization problem to avoid learning instability. In particular, the Proximal Policy Optimization (PPO) (Schulman et al., 2017) strategy, which we adopt in our framework, introduces constraints on policy updates through a clipped objective function. This ensures that updates do not destabilize the learning process by limiting the extent of policy changes, thereby providing stability and efficiency in fine-tuning large language models.
@@ 9       $j\leftarrow j+1$
  2. (2)
    Sentiment inference. We use a pre-trained model $\mathcal{S}$ to compute a sentiment value $s_{t}$ associated with the content $t$, i.e., $\mathcal{S}(t)=s_{t}$ (line 4). This utility model is a plug-and-play component that can be adapted to different tasks. In our experiments, we adopt a DistilBERT model (Sanh et al., 2019) pretrained for sentiment classification as our function $\mathcal{S}$.
  3. (3)
    Fluency score. To ensure that $t$ conveys proper semantics, we compute a readability value associated with its content. While perplexity is a common measure for evaluating the generation quality of LLMs (Gonen et al., 2023), it is sensitive to vocabulary size and sentence length. Studies (Miaschi et al., 2020) highlight its low correlation with readability. Martinc et al. (2021), study the adequacy of several readability scores and assess their performance on several datasets. For our purposes, we adopt as our fluency function $\mathcal{F}$ the Flesch–Kincaid (FK) (Kincaid et al., 1975) formula (line 4). This score quantifies the grade level required to understand an English statement: the higher the value, the higher its complexity and syntactic quality. It is a crucial component of the final reward $\mathcal{R}$, as it ensures the LLM generates meaningful and fluent statements.
  4. (4)
@@ 5. Experiments Setup
Synthetic data. To rigorously evaluate our proposed methodology across diverse initial conditions, we devise a synthetic data generator, based on random network model inspired by (Cinus et al., 2022), and capable of producing realistic networks while allowing to control the key parameters influencing information diffusion, such as homophily, modularity, and opinion distribution. The generator takes as input five parameters – homophily level $\eta$, network modularity $\mu$, shape parameters $\alpha$ and $\beta$ for opinion distribution, the number of nodes $N$ – and outputs a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, also assigning an opinion value $x_{u}\in[0,1]$ to each node $u\in\mathcal{V}$. The generation process starts by constructing the set of edges $\mathcal{E}$ and assigning nodes to non-overlapping communities $c:\mathcal{V}\rightarrow C$ using the LFR model (Lancichinetti et al., 2008), known for producing community sizes that adhere to a realistic power-law distribution. The modularity parameter $\mu$ is an input of the LFR model and controls the community segregation within the network. Each community $c_{i}\in C$ gets assigned an initial opinion drawn from a Beta distribution [^3] Beta distribution has been chosen for its flexibility to accommodate various distribution shapes such as uniform, unimodal, and bimodal. $o_{c_{i}}\sim Beta(\alpha,\beta)$. For each node $u\in\mathcal{V}$, its opinion $x_{u}$ is determined by a Bernoulli trial with probability $\eta$ (i.e., homophily). If successful, the node adopts the opinion of its community $x_{u}=o_{c(u)}$; otherwise, it independently samples an opinion from the same Beta distribution.
@@
Figure 2. Different opinion settings (from left to right): positive/negative/neutral/uniform/uniform. All networks except for the last one exhibit high modularity and high homophily (i.e., neighbors have similar opinions). The last graph is indeed an example of network with high modularity but low homophily (i.e., neighbors’ opinions are heterogeneous). Colors represent node opinions, whose values are denoted by the color bar.
@@
Figure 3. Networks with low modularity (left) and high modularity (right). The former presents high connectivity among the nodes, while the latter exhibits clustered communities.
@@
Figure 4. Visualization of the $\mathbb{X}$ follow networks. Nodes are colored on a scale from blue (0: ”Remain”/”No”) to red (1: ”Leave”/”Yes”), and their size resembles their out-degree.
@@
Real data. We also consider two real-world datasets containing content about a controversial topics propagated over $\mathbb{X}$ social network. The first dataset, Brexit, focuses on the remain-leave discourse before the 2016 UK Referendum on exiting the EU (Zhu et al., 2020). We use the version in (Minici et al., 2022) consisting of 7,589 users, 532,459 directed follow relationships, and 19,963 tweets, each associated with a binary stance. The dataset is preprocessed according to (Minici et al., 2022) to assign each user a scalar value $x_{u}\in[0,1]$, referred to as opinion, representing the average stance of the tweets retweeted by user $u$. The stance of each tweet is either 0 (“Remain”) or 1 (“Leave”). The second dataset, Italian Referendum, was gathered during the Italian constitutional referendum in 2016 (Lai et al., 2018) and processed according to (Minici et al., 2022). This dataset includes 2,894 users, 161,888 edges, and 41,001 tweets. Each user is associated with a scalar value $x_{u}\in[0,1]$ , corresponding to their stance on the referendum question, where values represent support for “No” or “Yes”. The structure of the networks and users’ opinions are depicted in Figure 4. Both datasets exhibit two separated, homophilic, and polarized communities, and a smaller, more heterophilic mixed-opinions community.
@@
Settings. Although other modeling choices are possible [^4] For instance, the Generation from Scratch task: i.e., “Generate a post about [TOPIC]”., in our experiments we focus on a Query Completion Task, as we found that Query Completion achieves faster convergence. We defer a detailed study on the effects of task specification to future work. Specifically, in the synthetic setting, we use the query ($q=$Cats are the most…”), while in the real experiments we adopt $q=$“[Brexit]/[2016 Italian constitutional referendum] is the most…”. Given the query $q$, we obtain $t$ by concatenating $q$ and the generated text $g=\textrm{LLM}_{\theta}(q)$, i.e., $t=q\mathbin{\|}g$. To compute the opinion value $s_{t}$, we adopt a DistilBERT model (Sanh et al., 2019), pretrained for sentiment classification.
@@
Concerning the LLM agent, we adopt the 2B version of Gemma (Team and et. al., 2024), a lightweight models family released by Google and based on Gemini [^5] https://deepmind.google/technologies/gemini/ technology. Despite its smaller size, Gemma outperforms other models of similar size in tasks such as understanding, reasoning, and safety. We fine-tune the model exploiting the PPOTrainer [^6] https://huggingface.co/docs/trl/main/en/ppo_trainer#trl.PPOTrainer class from the trl package, which enables training language models with custom rewards. The maximum number of training steps is set to $80$ for the synthetic data and $500$ for the real-world datasets, while the threshold $\tau$ for monitoring the KL-divergence to $75$. We also conducted experiments with other state-of-the-art lightweight LLMs, such as Mistral-7B (Jiang et al., 2023), LLaMA2-7B (Touvron et al., 2023), and GPT-2 (Radford et al., 2019). However, these experiments demonstrated that Gemma-2B provides the best trade-off between performance and computational efficiency. Finally, for the engagement model, we set $\epsilon=0.2$.
@@
Figure 5. Engagement $|\mathcal{A}|$ at each step of our fine-tuning procedure. Columns refer to different opinion distributions: positive/negative/neutral/uniform (left to right). Each plot depicts the trend varying the network structure in terms of modularity and homophily. Colors indicate different positions of the LLM agent. Dashed lines represent the maximum engagement within that configuration, whereas dotted lines indicate its lower-bound (i.e., $|\mathcal{A}|$ produced by the non-finetuned LLM).
@@
Figure 6. Sentiment content $s_{t}$ produced at each step of our fine-tuning procedure. Each column refers to a different opinion distribution: positive/negative/neutral/uniform (from left to right). Each plot depicts the trend varying the network structure in terms of modularity and homophily. Colors indicate different positions of the LLM agent.
Baselines. We evaluate the effectiveness of our approach to fine-tuning the LLM, comparing it with the following baselines: BERT [^7] https://huggingface.co/google-bert/ (Kenton and Toutanova, 2019), GPT-2 (Radford et al., 2019), LLaMA3.1-70B-Instruct [^8] https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct, and ChatGPT-4o [^9] https://openai.com/index/hello-gpt-4o/, in addition to the vanilla (not fine-tuned) Gemma-2B model (Team and et. al., 2024). We invoke each model in inference to complete the same query $q$ used in the synthetic experiments (“Cats are the most”) and in the real-world scenarios (“[Brexit]/[2016 Italian constitutional referendum] is the most”). When using BERT, we start from $q$ and masking the next word up to a final length of $10$. For all the other baselines, we set a maximum number of output tokens equal to $128$ and a temperature of $0.7$. Consistently with the fine-tuning process, we retrieve a number of responses equal to the batch size (8), and we further inject them into the network starting from the same source nodes. We average the results in terms of produced engagement.